A retail jewelry store has generated the following regression model for forecasting its total quarterly sales (in thousands of dollars):

^
Y3=81.5+4.08t-32.60Q1-16.30Q2+8.15Q3
This model is developed using a data set that starts in Quarter 4 of 2006 (i.e., the first time period t=1 is associated with Quarter 4 of 2006 ). Use this model to forecast the store's total sales (in thousands of dollars) in Quarter 2 of 2012.
I have t= 23
y23= 81.5+4.08(23)-32.6(0)-16.30(1)+8.15(0)=69.39 but is wrong do not know what I did wrong. can someone please help me

Answer by stanbon(75887)

Y3=81.5+4.08t-32.60Q1-16.30Q2+8.15Q3
This model is developed using a data set that starts in Quarter 4 of 2006 (i.e., the first time period t=1 is associated with Quarter 4 of 2006 ). Use this model to forecast the store's total sales (in thousands of dollars) in Quarter 2 of 2012.
I have t= 23
y23= 81.5+4.08(23)-32.6(0)-16.30(1)+8.15(0)=69.39 but is wrong do not know what I did wrong.
----------------
What do you mean by "Y3"?
Is that Y^3 or is it something else?
---
How did you determine that Q1 = 0,
and that Q2 = 1 and Q3 = 0?
=============
Using your numbers I get y23 = 81.5-16.3 = 65.2
Where did you get 69.39?
=============
Cheers,
Stan H.